Exercise 14.1. For each of the following functions F , determine if it is linear by using the viewpoint of behavior with respect to vector addition and scalar multiplication. If it is not linear, either find a pair of vectorsv,wforwhichF(v+w) ̸= F(v)+F(w)orfindavectorvandscalarcforwhichF(cv) ̸= cF(v).
Answer #1

htv-G). W= ) SR? Then f() = ((?:)) = 39] fCW) = f (05:)) - Cu toat] Now it w = {4, + ( + 1 = ( x + ( 2 ) - 343.470 (024-39-339h([7])=G) • Let u = w = ER? 2 2. Then, t(il= hlw= ūtw = 4th Sith 12, + z hlutw) = Z+22] nity I+Y =h(5)th(W) ..ni tw) = 4(5)t